# DRL-based Latency-Aware Network Slicing in O-RAN with Time-Varying SLAs

### v2

@@ Abstract
The Open Radio Access Network (Open RAN) paradigm, and its reference architecture proposed by the O-RAN Alliance, is paving the way toward open, interoperable, observable and truly intelligent cellular networks. Crucial to this evolution is Machine Learning (ML), which will play a pivotal role by providing the necessary tools to realize the vision of self-organizing O-RAN systems. However, to be actionable, ML algorithms need to demonstrate high reliability, effectiveness in delivering high performance, and the ability to adapt to varying network conditions, traffic demands and performance requirements. To address these challenges, in this paper we propose a novel Deep Reinforcement Learning (DRL) agent design for O-RAN applications that can learn control policies under varying Service Level Agreement (SLAs) with heterogeneous minimum performance requirements. We focus on the case of RAN slicing and SLAs specifying maximum tolerable end-to-end latency levels. We use the OpenRAN Gym open-source environment to train a DRL agent that can adapt to varying SLAs and compare it against the state-of-the-art. We show that our agent maintains a low SLA violation rate that is 8.3×8.3$\times$8.3 × and 14.4×14.4$\times$14.4 × lower than approaches based on Deep Q-Learning (DQN) and Q-Learning, while consuming respectively 0.3×0.3$\times$0.3 × and 0.6×0.6$\times$0.6 × less resources without the need for re-training.
Index Terms:
O-RAN, Network Slicing, Deep Reinforcement Learning, Latency-aware, Service Level Agreement.
@@ I Introduction
To effectively cater to a wide spectrum of application-specific and user-centric demands, there is a need to rethink the way Radio Access Networks (RANs) are designed and deployed. Industry and academia alike commonly agree that this can be achieved by combining three core principles [1]: (i) programmable and virtualized protocol stacks with clearly defined open interfaces; (ii) closed-loop and fast network control and reconfiguration; and (iii) data-driven modeling and machine learning (ML). These constitute the foundational principles of the emergent Open RAN paradigm, which has recently garnered substantial momentum as a practical enabler of algorithmic and hardware innovation in future cellular networks [1, 2]. The O-RAN Alliance [3] is establishing a reference architecture for the Open RAN and defining its specifications with the ultimate goal to push Open RAN toward standardization. In this context, the O-RAN Alliance is promoting open interfaces that connect the various disaggregated functional units, and enable their programmability and monitoring by a set of RAN Intelligent Controllers (RICs). Specifically, O-RAN introduces the Non-real-time (Non-RT) the Near-RT RICs. The former hosts intelligent algorithms called rApps that perform inference at timescales larger than $1$s, while the latter hosts xApps that operate at timescales from $10$ms to $1$s. Intelligent and dynamic network optimization via xApps and rApps introduces novel practical challenges, which include, on one side, designing machine learning (ML) agents capable of adapting to unseen conditions and deployments, and on the other side, selecting features that provide a meaningful representation of the network status without incurring in dimensionality explosion [4]. Moreover, advances in ML-based network automation have been slow, mainly due to the unavailability of large-scale datasets and experimental testing infrastructure to train, test and validate xApps and rApps at scale [5].
@@
To address these key challenges, in this paper, we introduce a novel design of DRL-based xApps for closed-loop control in O-RAN with the objective of computing slicing policies that can satisfy varying SLAs for different network slices while minimizing the amount of Physical Resource Blocks (PRBs) allocated to mobile users, so as to improve resource utilization efficiency and reduce energy consumption. We discuss practical challenges of integrating latency-aware DRL algorithms within O-RAN (e.g., how to generate latency measurements and which interfaces to use to make them available to the RIC), and show that our solution can adapt to varying SLAs while improving upon DQN and Q-learning approaches in the literature. Specifically, we show that we can reduce SLA violations by 8.3×8.3$\times$8.3 × and 14.4×14.4$\times$14.4 × while consuming 0.3×0.3$\times$0.3 × to 0.6×0.6$\times$0.6 × less PRBs.
@@ III-A O-RAN architecture overview
An important O-RAN innovation is the RIC which offers an abstraction and virtualization layer to support various control loops operating at different timescales, ranging from near-real-time ($10$ms to $1$s) to non-real-time (higher than $1$s). The Near-RT RIC hosts xApps that receive data and telemetry over the E2 interface from the so-called E2 nodes (i.e., DUs, CUs). The E2 interface is logically structured into two distinct protocols: the E2 Application Protocol (E2AP) and E2 Service Model (E2SM). E2AP serves as a procedural protocol, coordinating communication between the near-RT RIC and E2 nodes, while E2SMs, embedded in E2AP messages, implement specific functionalities, such as reporting RAN metrics or controlling RAN parameters. The Non-RT RIC hosts rApps which, similarly to xApps, embed intelligent algorithms to monitor and control the RAN. The Non-RT RIC collects data and control RAN components via the O1 interface, and is connected via the A1 interface to the Near-RT RIC. The A1 interface is used by the Non-RT RIC to exchange enrichment information and policies with the Near-RT RIC. The Non-RT RIC is then embedded within the Service Management and Orchestration (SMO) component which handles lifecycle of all O-RAN components.
@@ V-A STAT Scenario: Performance evaluation and comparison
Regarding the second network slice, and as shown in Figs. 6 and 6, we notice that the Q-Learning approach still fails to yield a satisfactory solution. DRL-based solutions exhibit better performance if compared to Q-Learning, but consistently fail in satisfying SLAs. Our proposed approach stands out as the sole method capable of fulfilling these requirements. This distinction is further corroborated by the episode reward, which averages 4.5$\times$ and 12$\times$ higher than the DQN-based and Q-learning-based solutions, respectively.
@@ V-B DYN Scenario: performance evaluation and comparison
In Figs. 8 and 8, we show the reward and violation rate of different agents during training in the DYN Scenario. Fig. 8 shows that our proposed approach outperforms the DQN-based and Q-Learning approaches by delivering a higher reward. Similarly, Fig. 8 shows that our approach also delivers the lowers violation rate. This is, on average, 8.3×8.3$\times$8.3 × and 14.4×14.4$\times$14.4 × lower than DQN and Q-Learning, respectively.
@@ VI Conclusions and Future Work
In this paper, we have presented the design of a DRL agent that can learn to react to varying SLAs without requiring any retraining. We design the DRL agent in a way that makes it suitable for O-RAN applications and illustrate the integration efforts required to execute our agent as an xApp hosted in the Near-RT RIC. We train our agent using a PPO architecture on data collected using OpenRAN Gym and Colosseum, and show that our solution can adapt to varying SLAs while improving upon DQN and Q-learning approaches in the literature. Specifically, we show that we can reduce SLA violations by 8.3×8.3$\times$8.3 × and 14.4×14.4$\times$14.4 × while consuming from 0.3×0.3$\times$0.3 × to 0.6×0.6$\times$0.6 × less PRBs. In the future, we will investigate the case of multi-agent DRL solutions coordinating decisions to simultaneously satisfy diverse SLAs, with possibly different target KPMs.

### v3

@@ Abstract
The Open Radio Access Network (Open RAN) paradigm, and its reference architecture proposed by the O-RAN Alliance, is paving the way toward open, interoperable, observable and truly intelligent cellular networks. Crucial to this evolution is Machine Learning (ML), which will play a pivotal role by providing the necessary tools to realize the vision of self-organizing O-RAN systems. However, to be actionable, ML algorithms need to demonstrate high reliability, effectiveness in delivering high performance, and the ability to adapt to varying network conditions, traffic demands and performance requirements. To address these challenges, in this paper we propose a novel Deep Reinforcement Learning (DRL) agent design for O-RAN applications that can learn control policies under varying Service Level Agreement (SLAs) with heterogeneous minimum performance requirements. We focus on the case of RAN slicing and SLAs specifying maximum tolerable end-to-end latency levels. We use the OpenRAN Gym open-source environment to train a DRL agent that can adapt to varying SLAs and compare it against the state-of-the-art. We show that our agent maintains a low SLA violation rate that is $8.3\times$ and $14.4\times$ lower than approaches based on Deep Q-Learning (DQN) and Q-Learning, while consuming respectively $0.3\times$ and $0.6\times$ less resources without the need for re-training.
@@ I Introduction
To effectively cater to a wide spectrum of application-specific and user-centric demands, there is a need to rethink the way Radio Access Networks (RANs) are designed and deployed. Industry and academia alike commonly agree that this can be achieved by combining three core principles [1]: (i) programmable and virtualized protocol stacks with clearly defined open interfaces; (ii) closed-loop and fast network control and reconfiguration; and (iii) data-driven modeling and machine learning (ML). These constitute the foundational principles of the emergent Open RAN paradigm, which has recently garnered substantial momentum as a practical enabler of algorithmic and hardware innovation in future cellular networks [1, 2]. The O-RAN Alliance [3] is establishing a reference architecture for the Open RAN and defining its specifications with the ultimate goal to push Open RAN toward standardization. In this context, the O-RAN Alliance is promoting open interfaces that connect the various disaggregated functional units, and enable their programmability and monitoring by a set of RAN Intelligent Controllers (RICs). Specifically, O-RAN introduces the Non-real-time (Non-RT) the Near-RT RICs. The former hosts intelligent algorithms called rApps that perform inference at timescales larger than $1$s, while the latter hosts xApps that operate at timescales from 1010ms to $1$s. Intelligent and dynamic network optimization via xApps and rApps introduces novel practical challenges, which include, on one side, designing machine learning (ML) agents capable of adapting to unseen conditions and deployments, and on the other side, selecting features that provide a meaningful representation of the network status without incurring in dimensionality explosion [4]. Moreover, advances in ML-based network automation have been slow, mainly due to the unavailability of large-scale datasets and experimental testing infrastructure to train, test and validate xApps and rApps at scale [5].
@@
To address these key challenges, in this paper, we introduce a novel design of DRL-based xApps for closed-loop control in O-RAN with the objective of computing slicing policies that can satisfy varying SLAs for different network slices while minimizing the amount of Physical Resource Blocks (PRBs) allocated to mobile users, so as to improve resource utilization efficiency and reduce energy consumption. We discuss practical challenges of integrating latency-aware DRL algorithms within O-RAN (e.g., how to generate latency measurements and which interfaces to use to make them available to the RIC), and show that our solution can adapt to varying SLAs while improving upon DQN and Q-learning approaches in the literature. Specifically, we show that we can reduce SLA violations by $8.3\times$ and $14.4\times$ while consuming $0.3\times$ to $0.6\times$ less PRBs.
@@ III-A O-RAN architecture overview
An important O-RAN innovation is the RIC which offers an abstraction and virtualization layer to support various control loops operating at different timescales, ranging from near-real-time (1010ms to $1$s) to non-real-time (higher than $1$s). The Near-RT RIC hosts xApps that receive data and telemetry over the E2 interface from the so-called E2 nodes (i.e., DUs, CUs). The E2 interface is logically structured into two distinct protocols: the E2 Application Protocol (E2AP) and E2 Service Model (E2SM). E2AP serves as a procedural protocol, coordinating communication between the near-RT RIC and E2 nodes, while E2SMs, embedded in E2AP messages, implement specific functionalities, such as reporting RAN metrics or controlling RAN parameters. The Non-RT RIC hosts rApps which, similarly to xApps, embed intelligent algorithms to monitor and control the RAN. The Non-RT RIC collects data and control RAN components via the O1 interface, and is connected via the A1 interface to the Near-RT RIC. The A1 interface is used by the Non-RT RIC to exchange enrichment information and policies with the Near-RT RIC. The Non-RT RIC is then embedded within the Service Management and Orchestration (SMO) component which handles lifecycle of all O-RAN components.
@@ V-A STAT Scenario: Performance evaluation and comparison
[…]
@@ V-B DYN Scenario: performance evaluation and comparison
In Figs. 8 and 8, we show the reward and violation rate of different agents during training in the DYN Scenario. Fig. 8 shows that our proposed approach outperforms the DQN-based and Q-Learning approaches by delivering a higher reward. Similarly, Fig. 8 shows that our approach also delivers the lowers violation rate. This is, on average, $8.3\times$ and $14.4\times$ lower than DQN and Q-Learning, respectively.
@@ VI Conclusions and Future Work
In this paper, we have presented the design of a DRL agent that can learn to react to varying SLAs without requiring any retraining. We design the DRL agent in a way that makes it suitable for O-RAN applications and illustrate the integration efforts required to execute our agent as an xApp hosted in the Near-RT RIC. We train our agent using a PPO architecture on data collected using OpenRAN Gym and Colosseum, and show that our solution can adapt to varying SLAs while improving upon DQN and Q-learning approaches in the literature. Specifically, we show that we can reduce SLA violations by $8.3\times$ and $14.4\times$ while consuming from $0.3\times$ to $0.6\times$ less PRBs. In the future, we will investigate the case of multi-agent DRL solutions coordinating decisions to simultaneously satisfy diverse SLAs, with possibly different target KPMs.